# Understanding the Drag Torque in Common Envelope Evolution

### v1

@@ 5.2.3 Limitations that are common to the simulation and K08 model
Thus far, we have discussed the limitations of the K08 model for reproducing our simulation results, but the K08 model and our CE simulations also have certain assumptions in common that are not fully realistic. First, the equation of state is taken to be that of an adiabatic ideal gas with $\gamma=5/3$, which means that ionization, recombination, radiative cooling, and radiative transfer are neglected. Both hydrogen and helium are expected to be completely ionized near the perturbers, the envelope is optically thick, and the orbital evolution is fairly insensitive to the equation of state (e.g. Sand et al., 2020; Chamandy et al., 2024). Thus, we do not expect that the torque evolution will depend very significantly on the equation of state and the inclusion of radiation during the time interval considered. However, at very late times not reached by our simulation the gas around the perturbers would become optically thin, at which point these assumptions would be invalid and the modeled torque inaccurate.
@@
Another assumption shared by K08 and our simulation is that the perturbers do not accrete or launch jets/winds. The accretion rate during CE evolution might, in some cases, attain hyper-critical values orders of magnitude larger than the Eddington rate, in which case it could affect significantly the orbital separation through its effect on the gravitational torque (Chamandy et al., 2018). However, this effect is unlikely to be important in cases where feedback prevents the accretion rate from greatly exceeding the Eddington rate. Relevant work on dynamical friction drag on accreting perturbers in a gaseous medium is presented in Ruffert (1996), Sánchez-Salcedo & Brandenburg (2001), Gruzinov et al. (2020), Prust et al. (2024), and Suzuguchi et al. (2024). A somewhat related issue is the lack of sophistication in the modeling of the AGB core and companion, which are represented by point particles. While there has been recent work on modeling the perturbers as spheres (Prust, 2020; Prust & Bildsten, 2024), this may not make a major difference since the mass distribution around the perturbers tends to be quasi-hydrostatic in the CE simulations (e.g. Kim 2010, Chamandy et al. 2018; see also the top row of Figure 7), mimicking, in a loose sense, the structure of a star.
@@
Other effects that are not modeled are turbulence (possibly convective) and MHD effects. Actually, turbulence seems to be captured to some extent in our CE simulations (Chamandy et al. 2019a; see also second row of Figure 7, where patchy morphology suggests the presence of turbulence), but turbulence is not considered in K08, though it may have a significant effect on the drag (Lescaudron et al., 2023). Magnetohydrodynamics (MHD) is not considered in any of the above models, though magnetic fields can affect the dynamical friction force (Sánchez-Salcedo, 2012; Shadmehri & Khajenabi, 2012). However, CE simulation studies have found that magnetic fields only weakly affect the orbital evolution, likely because the magnetic fields that arise in the simulations are themselves fairly weak near the particles (Ohlmann et al., 2016b; Ondratschek et al., 2022). However, Ondratschek et al. (2022) and Vetter et al. (2024) find that the magnetic fields are strong enough to produce powerful bipolar outflows. More studies are needed to better understand the role of magnetic fields in CE evolution.
@@ 6 SUMMARY AND CONCLUSIONS
  • We first interpreted our results using a model akin to that of E04, modeling the torque on the particles as that produced by a uniform density ellipsoid rotating with the same angular velocity as the particles but lagging by a phase angle. The simplest version of this model invokes a prolate spheroid with constant axis ratio and lag angle. These models reproduce remarkably well the torque as a function of time, but contain parameters, viz. the major axis to particle separation ratio (or, equivalently, the mean density), spheroid axis ratio, and lag angle, that needed to be estimated from the simulation.
@@
  • The best-fit lag angle, axis ratio, and semi-major axis to separation ratio are found to be remarkably close to those found by E04 (their figure 8), who simulated BSMBHs; it is unclear to what extent this points to universal behaviour in different astrophysics contexts or is more of a coincidence.
  • K08 developed a different, first principles model to explain the torque on double perturber systems, and used it to explain the results of E04; see the bottom row of Figure 3 for an illustration of the consistency between the two types of model. We applied the K08 model to our simulation and again found remarkable agreement with the measured torque magnitude, amplitude, and the temporal phase of the torque oscillations caused by orbital eccentricity. This level of agreement was obtained by setting the unperturbed density, a parameter of the model, to $0.44$ times the density at the surface of the torque-dominating region around the particles, and by setting the perturber Mach number to the ratio of the particle tangential speed to the mean sound speed in the torque-dominating region (bottom panel of Figure 2).
  • This factor of $0.44$, an adjustable parameter of our model, is not needed to reproduce the magnitude of the torque at very late times ($t>270\,{\rm d}$), but we speculate that the torque at the very end of the simulation may be too high because the separation to softening radius ratio has become quite small by then.
  • Using instead the density and sound speed of the initial AGB profile to predict the unperturbed density and sound speed parameter in the K08 model does not result in good agreement. But perhaps a more complicated relationship between the initial profile and model parameters still exists, and this possibility should be explored in future work.
  • We argued in Section 5.2 that while there are many extra effects in the CE phase that are not included in the K08 model, e.g. gas self-gravity, nonlinear density perturbations, eccentricity, and perturber CM motion, these effects affect the torque only weakly. Hence the high level of agreement observed between the K08 idealized model and our CE simulation. The underlying reason for this may be that the flow adjusts on the sound-crossing timescale, which is shorter than the other timescales in the problem, e.g. $t_{\mathrm{s}}<P_{\mathrm{orb}}<t_{\mathrm{ff}}<a/\dot{a}$.
  • That nonlinear effects, in particular, seem to be unimportant may be a consequence of the perturbers’ motion being subsonic (Kim & Kim, 2009).
@@
Further work is needed to learn how to interpret the torque in other CE simulations in terms of the K08 model or future improvements thereof, and constrain the model parameters. Our results strongly suggest that numerous simulations showing a long-term secular decrease in the binary separation were, in fact, correct. They also hint at how one could eventually predict, perhaps with some accuracy, the timescale for this orbital tightening to occur. Down the road, this may lead to an accurate assessment of CE outcomes (e.g., merger or envelope ejection) and, more generally, to simple time-dependent models of CE evolution.
@@
We thank Andrés Escala for words of encouragement and Abha Vishwakarma for discussions.This work used the computational and visualization resources in the Center for Integrated Research Computing (CIRC) at the University of Rochester and the computational resources of the Rosen Center for Advanced Computing (RCAC) at Purdue University, provided through allocations PHY230168 and TG-TRA210040 from the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) program (Boerner et al., 2023), which is supported by National Science Foundation grants #2138259, #2138286, #2138307, #2137603, and #2138296, and the computational resources on Frontera of the Texas Advanced Computing Center (TACC) at The University of Texas at Austin, provided through TACC Frontera Pathways allocation AST23022. Frontera (Stanzione et al., 2020) is made possible by National Science Foundation award OAC-1818253.
@@
$&\vec{F}=-M\nabla\Phi=-\pi G\overline{\rho}M\\ &\times\bigg{\{}\left[2x\left(\alpha\cos^{2}(\Delta\phi)+\beta\sin^{2}(\Delta\phi)\right)+y\sin(2\Delta\phi)(\beta-\alpha)\right]{{\mbox{\boldmath$\hat{x}$}}}\\ &+\left[x\sin(2\Delta\phi)(\beta-\alpha)+2y\left(\alpha\sin^{2}(\Delta\phi)+\beta\cos^{2}(\Delta\phi)\right)\right]{{\mbox{\boldmath$\hat{y}$}}}\\ &+2\zeta z\,{{\mbox{\boldmath$\hat{z}$}}}\bigg{\}}.\end{aligned}$ (29)
@@
$&\tau=\pi G\overline{\rho}M\\ &\times\bigg{\{}z\left[x\sin(2\Delta\phi)(\beta-\alpha)\right.\\ &\left.+2y\left(\alpha\sin^{2}(\Delta\phi)+\beta\cos^{2}(\Delta\phi)-\zeta\right)\right]{{\mbox{\boldmath$\hat{x}$}}}\\ &-z\left[2x\left(\alpha\cos^{2}(\Delta\phi)+\beta\sin^{2}(\Delta\phi)+\zeta\right)\right.\\ &\left.+y\sin(2\Delta\phi)(\beta-\alpha)\right]{{\mbox{\boldmath$\hat{y}$}}}\\ &+\left[\sin(2\Delta\phi)(\beta-\alpha)(y^{2}-x^{2})\right.\\ &\left.+2xy\Big{(}\alpha\cos(2\Delta\phi)-\beta\cos(2\Delta\phi)\Big{)}\right]{{\mbox{\boldmath$\hat{z}$}}}\bigg{\}}\end{aligned}$ (30)
@@
Appendix B Evolution of gas density, sound speed, and particle velocities
@@
In Figure 9, we present the evolution of the magnitudes and $\phi$-components of the particle velocities (which are almost equal since the azimuthal orbital component dominates). Also shown are the sound speed averaged inside the $\rho=\rho_{\mathrm{c}}$ contour, on the contour, or as obtained from the original profile by setting $r=a(t)$, where $a$ is the particle separation.
@@
Figure 9: Velocity magnitude of the AGB core particle (red) and companion (blue), along with the $\phi$-components of the particle velocities (dark red and cyan, for the core and companion, respectively). The $\phi$-components can be seen to coincide very closely with the total speeds. The green lines show the adiabatic sound speed; dashed, solid and dotted lines respectively show the mean value inside the surface $\rho=\rho_{\mathrm{c}}$, the value on this surface, and the value at $r=a(t)$ in the original profile. The orange lines show the Mach number ($v_{\phi}/\overline{c}_{\mathrm{s}}$) of the AGB core particle (solid) and companion (dashed). The total torque on the binary (black) is plotted on the right axis for reference.
Appendix C Energy conservation
@@
Appendix D Envelope unbinding
In Figure 11 we present the evolution of the unbound and total gas mass, integrated over the simulation domain. Solid lines take into account the flux through the boundaries, but dashed lines do not. One can see that total gas mass is conserved in the simulation (and the total mass, since the particle masses are kept constant). For a comparison with a different simulation that is the same in all respects to the one studied in this work (including the same AGB primary star, initial separation, and initialization in a circular orbit) except that the companion is more massive ($0.98\,\mathrm{M_{\odot}}$ instead of $0.53\,\mathrm{M_{\odot}}$), we refer the reader to figure 7 of Chamandy et al. (2020). The simulations have approximately the same duration but the $0.98\,\mathrm{M_{\odot}}$ companion simulation inspirals to a somewhat larger final separation and unbinds somewhat more mass than the $0.53\,\mathrm{M_{\odot}}$ simulation presented in this work.

### v3

@@ 5.2.3 Limitations that are common to the simulation and Kim+08 model
Thus far, we have discussed the limitations of the Kim+08 model for reproducing our simulation results, but the Kim+08 model and our CE simulations also have certain assumptions in common that are not fully realistic. First, the equation of state is taken to be that of an adiabatic ideal gas with $\gamma=5/3$, which means that ionization, recombination, radiative cooling, and radiative transfer are neglected. Both hydrogen and helium are expected to be completely ionized near the perturbers, the envelope is optically thick, and the orbital evolution is fairly insensitive to the equation of state (e.g. Sand+20; Chamandy+24). Thus, we do not expect that the torque evolution will depend very significantly on the equation of state and the inclusion of radiation during the time interval considered. However, at very late times not reached by our simulation the gas around the perturbers would become optically thin, at which point these assumptions would be invalid and the modeled torque inaccurate.
@@
Another assumption shared by Kim+08 and our simulation is that the perturbers do not accrete or launch jets/winds. The accretion rate during CE evolution might, in some cases, attain hyper-critical values orders of magnitude larger than the Eddington rate, in which case it could affect significantly the orbital separation through its effect on the gravitational torque (Chamandy+18). However, this effect is unlikely to be important in cases where feedback prevents the accretion rate from greatly exceeding the Eddington rate. Relevant work on dynamical friction drag on accreting perturbers in a gaseous medium is presented in Ruffert96, Sanchez-salcedo+Brandenburg01, Gruzinov+20, Prust+24, and Suzuguchi+24. A somewhat related issue is the lack of sophistication in the modeling of the AGB core and companion, which are represented by point particles. While there has been recent work on modeling the perturbers as spheres (Prust20; Prust+Bildsten24), this may not make a major difference since the mass distribution around the perturbers tends to be quasi-hydrostatic in the CE simulations (e.g. Kim10, Chamandy+18; see also the top row of Figure 7), mimicking, in a loose sense, the structure of a star.
@@
Other effects that are not modeled are turbulence (possibly convective) and MHD effects. Actually, turbulence seems to be captured to some extent in our CE simulations (Chamandy+19b; see also second row of Figure 7, where patchy morphology suggests the presence of turbulence), but turbulence is not considered in Kim+08, though it may have a significant effect on the drag (Lescaudron+23). Magnetohydrodynamics (MHD) is not considered in any of the above models, though magnetic fields can affect the dynamical friction force (Sanchez-salcedo12; Shadmehri+Khajenabi12). While certain CE simulation studies have found that magnetic fields only weakly affect the orbital evolution (Ohlmann+16b; Ondratschek+22), Vetter+25 find that the inspiral is significantly deeper at late times in their MHD run as compared to their pure hydrodynamics run. Moreover, Ondratschek+22, Vetter+24 and Vetter+25 find that the magnetic fields are strong enough to produce powerful bipolar outflows. More studies are needed to better understand the role of magnetic fields in CE evolution.
5.3 Non-applicability to Very Early or Very Late Times
The duration of our simulation is about $40$ orbits and the models are applied for the final $36$ orbits. The symmetry that makes the Escala+04 and Kim+08 models applicable over the last roughly $36$ orbits is not present during the initial plunge-in. At those times models based on wind-tunnel simulations (e.g Macleod+Ramirez-ruiz15a; Macleod+17; De+20), or single perturbers (e.g. Kim+Kim07; Kim10) may be effective. At very late times (perhaps $\gtrsim 10^{3}$ orbits) the envelope will be largely ejected, and self-similar evolution is unlikely to persist. Some CE simulations reveal an asymmetric, partly evacuated region in the orbital plane just outside the binary at these late times (Gagnier+Pejcha25; Vetter+25), so the radial density profile can be non-monotonic. Different torque models may be required to explain these late stages of evolution.
@@ 6 SUMMARY AND CONCLUSIONS
  • We first interpreted our results using a model akin to that of Escala+04, modeling the torque on the particles as that produced by a uniform density ellipsoid rotating with the same angular velocity as the particles but lagging by a phase angle. The simplest version of this model invokes a prolate spheroid with constant axis ratio and lag angle. These models reproduce remarkably well the torque as a function of time, but contain parameters, viz. the major axis to particle separation ratio (or, equivalently, the mean density), spheroid axis ratio, and lag angle, that needed to be estimated from the simulation.
@@
  • The best-fit lag angle, axis ratio, and semi-major axis to separation ratio are found to be remarkably close to those found by Escala+04 (their figure 8), who simulated BSMBHs; it is unclear to what extent this points to universal behaviour in different astrophysics contexts or is more of a coincidence.
  • Kim+08 developed a different, first principles model to explain the torque on double perturber systems, and used it to explain the results of Escala+04; see the bottom row of Figure 3 for an illustration of the consistency between the two types of model. We applied the Kim+08 model to our simulation and again found remarkable agreement with the measured torque magnitude, amplitude, and the temporal phase of the torque oscillations caused by orbital eccentricity. This level of agreement was obtained by setting the unperturbed density, a parameter of the model, to $\xi=0.44$ times the density at the surface of the torque-dominating region around the particles, and by setting the perturber Mach number to the ratio of the particle tangential speed to the mean sound speed in the torque-dominating region (bottom panel of Figure 2). The best-fit value of $\xi$ may be different for other common envelope systems with different initial conditions.
  • While our model reproduces the magnitude of the torque at very late times ($t>270\,{\rm d}$) with $\xi$ set to unity, this is likely a coincidence: the torque at the very end of the simulation may be overestimated because the separation to softening radius ratio has become quite small by then.
  • Using instead the density and sound speed of the initial AGB profile to predict the unperturbed density and sound speed parameter in the Kim+08 model does not result in good agreement. But perhaps a more complicated relationship between the initial profile and model parameters still exists, and this possibility should be explored in future work.
  • We argued in Section 5.2 that while there are many extra effects in the CE phase that are not included in the Kim+08 model, e.g. gas self-gravity, nonlinear density perturbations, eccentricity, and perturber CM motion, these effects affect the torque only weakly. Hence the high level of agreement observed between the Kim+08 idealized model and our CE simulation. The underlying reason for this may be that the flow adjusts on the sound-crossing timescale, which is shorter than the other timescales in the problem, e.g. $t_{\mathrm{s}}<P_{\mathrm{orb}}<t_{\mathrm{ff}}<a/\dot{a}$.
  • That nonlinear effects, in particular, seem to be unimportant may be a consequence of the perturbers’ motion being subsonic (Kim+Kim09).
@@
Further work is needed to learn how to interpret the torque in other CE simulations in terms of the Kim+08 model or future improvements thereof, and constrain the model parameters. Our results strongly suggest that numerous simulations showing a long-term secular decrease in the binary separation were, in fact, correct. They also hint at how one could eventually predict, perhaps with some accuracy, the timescale for this orbital tightening to occur. Down the road, this may lead to an accurate assessment of CE outcomes (e.g., merger or envelope ejection) and, more generally, to increasingly accurate simple time-dependent models of CE evolution (e.g., Fragos+19; Bronner+24).
@@
We are grateful to the reviewer for a careful reading of the manuscript and helpful suggestions that improved the work. We thank Andrés Escala for words of encouragement and Shubhajit Saha and Abha Vishwakarma for discussions. This work used the computational and visualization resources in the Center for Integrated Research Computing (CIRC) at the University of Rochester and the computational resources of the Rosen Center for Advanced Computing (RCAC) at Purdue University, provided through allocations PHY230168 and TG-TRA210040 from the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) program (Access), which is supported by National Science Foundation grants #2138259, #2138286, #2138307, #2137603, and #2138296, and the computational resources on Frontera of the Texas Advanced Computing Center (TACC) at The University of Texas at Austin, provided through TACC Frontera Pathways allocation AST23022. Frontera (Frontera) is made possible by National Science Foundation award OAC-1818253.
@@
[…]
@@
$\begin{aligned} &\tau=\pi G\overline{\rho}M\\ &\times\bigg\{z\left[x\sin(2\Delta\phi)(\beta-\alpha)\right.\\ &\left.+2y\left(\alpha\sin^{2}(\Delta\phi)+\beta\cos^{2}(\Delta\phi)-\zeta\right)\right]{{\mbox{$\hat{x}$}}}\\ &-z\left[2x\left(\alpha\cos^{2}(\Delta\phi)+\beta\sin^{2}(\Delta\phi)+\zeta\right)\right.\\ &\left.+y\sin(2\Delta\phi)(\beta-\alpha)\right]{{\mbox{$\hat{y}$}}}\\ &+\left[\sin(2\Delta\phi)(\beta-\alpha)(y^{2}-x^{2})\right.\\ &\left.+2xy\Big(\alpha\cos(2\Delta\phi)-\beta\cos(2\Delta\phi)\Big)\right]{{\mbox{$\hat{z}$}}}\bigg\}\end{aligned}$ (30)
@@
Appendix B Determination of $\rho_{\mathrm{c}}$
The critical density $\rho_{\mathrm{c}}$ defines a geometric surface that encloses the gas which approximately contributes all of the torque on the particles. We also require this surface to be roughly ellipsoidal in shape and small enough to be usefully applied to the torque models considered. We model $\rho_{\mathrm{c}}$ as a fraction $f$ of the maximum density $\rho_{\mathrm{max}}$ (situated near the location of the AGB core particle for all snapshots), and choose the value of $f$ which leads to a torque that is closest to the total torque exerted by all the gas in the simulation. In Fig. 8, we plot the total torque (black dash-dotted) and the torque for various values of $f$, as a function of time. On the right axis, we plot the percent deviation. The value $f=0.006$, shown in magenta, produces a torque closest to the total torque, with an average percent absolute deviation of $10.6\%$, as compared to $13.9\%$ for $0.007$ and $12.3\%$ for $0.005$ in the range $t=125\,{\rm d}$ to $t=300\,{\rm d}$. If we cap the range at $270\,{\rm d}$ to exclude the times when $r_{\mathrm{soft}}/a$ is likely too large, $0.006$ still gives the smallest deviation ($11.1\%$). Hence, we fix $f=0.006$ for this study.
Figure 8: Comparison of the torque on the particles contributed by gas inside the contour $\rho=\rho_{\mathrm{c}}=f\rho_{\mathrm{max}}$, with $\rho_{\mathrm{max}}$ the maximum density in the simulation, to the total torque contributed by all gas (dash-dotted black). The percent deviation from the total torque is shown on the right axis. For the torque models in this work we choose $f=0.006$, which gives the lowest time-averaged percent deviation of the values plotted for $t>125\,{\rm d}$.
Appendix C Evolution of gas density, sound speed, and particle velocities
@@
In Figure 10, we present the evolution of the magnitude and $\phi$-component of the velocity of each particle in the particle CM frame, $V_{1}(=V_{2})$ and $V_{1,\phi}(=V_{2,\phi})$, are seen to be almost equal since the azimuthal orbital component dominates. Also shown are the sound speed averaged inside the $\rho=\rho_{\mathrm{c}}$ contour, on the contour, or as obtained from the original profile by setting $r=a(t)$, where $a$ is the particle separation.
@@
Figure 10: Speed of the particles in the particle CM frame (solid blue), along with the $\phi$-component of the particle velocity (dotted cyan) (the speeds of the particles are equal by definition in this frame). The magnitude of the $\phi$-component can be seen to coincide very closely with the total speed. The green lines show the adiabatic sound speed; dashed, solid and dotted lines respectively show the mean value inside the surface $\rho=\rho_{\mathrm{c}}$, the value on this surface, and the value at $r=a(t)$ in the original profile. The solid orange line shows the Mach number ($V_{\phi}/\overline{c}_{\mathrm{s}}$) of the AGB core particle and companion, in the particle CM reference frame. The total torque on the binary (black) is plotted on the right axis for reference.
Appendix D Energy conservation
@@
Appendix E Envelope unbinding
In Figure 12 we present the evolution of the unbound and total gas mass, integrated over the simulation domain. Solid lines take into account the flux through the boundaries, but dashed lines do not. One can see that total gas mass is conserved in the simulation (and the total mass, since the particle masses are kept constant). For a comparison with a different simulation that is the same in all respects to the one studied in this work (including the same AGB primary star, initial separation, and initialization in a circular orbit) except that the companion is more massive ($0.98\,\mathrm{M_{\odot}}$ instead of $0.53\,\mathrm{M_{\odot}}$), we refer the reader to figure 7 of Chamandy+20. The simulations have approximately the same duration but the $0.98\,\mathrm{M_{\odot}}$ companion simulation inspirals to a somewhat larger final separation and unbinds somewhat more mass than the $0.53\,\mathrm{M_{\odot}}$ simulation presented in this work.